Subject: APO 3.5m users committee minutes, Oct 2014

From: Suzanne Hawley

Submitted: Thu, 6 Nov 2014 23:52:02 -0800 (PST)

Message number: 1346 (previous: 1345, next: 1347 up: Index)

 		APO 3.5-m Users Committee Phonecon, 10/27/14

Attending:  Suzanne Hawley, Mark Klaene, Rachel Kuzio de Naray, Scott 
Anderson, John Wisniewski, Michael Strauss, John Bally, Rene Walterbos, 
Mary Beth Kaiser, Remy Indebetouw, and Bruce Gillespie


User feedback and comments from institutional representatives:

  o Washington - Scott had no issues or concern.  For the 3.5-m/APOGEE 
fiber link, one person was possibly interested.

  o Virginia - Remy said there were no user complaints from UVa.

  o Georgia State - Rachel reported that everything is going well.  Nobody 
at GSU had an opinion on putting in a fiber link from the 3.5-m telescope 
to the APOGEE instrument.

  o Princeton - Michael noted that PU users were relatively quiet. 
Suzanne mentioned that there was a recent Princeton ToO request that was 

  o Colorado - John had no user concerns.  There is strong enthusiasm for a 
3.5-m/APOGEE fiber connection at Boulder.

  o Johns Hopkins - Mary Beth has heard no feedback from JHU users.

  o NMSU - No report.

  o Seattle - No report.

  o Oklahoma - John said that the Oklahoma users had no concerns.

  o NAPG - No report.


Discussion of telescope/instruments report:

Mark highlighted the unusually bad September weather; October has been 
better.  After 24 years of service, Gretchen Van Doren is retiring at the 
end of the year.  Her replacement, Beth Mitchell, will start in November. 
The telescope has been generally well-behaved, with only a few small 
problems related to the recent upgrades to the Telescope Control Computer 
hardware and software.  For the instruments, there are no big issues at 
the moment.  For example, the echelle inter-order stray light appears to 
be back to normal after warming and pumping the detector dewar. 
Engineers from the GSFC made improvements to the cryogenic systems on the 
GIFS that have significantly improved the LN2 hold time.  The CCD detector 
for the new ARCTIC imager was not delivered as promised and we are 
discussing options.  Work continues on the optics and electronics for 
ARCTIC.  Bruce noted that there were some recent reports of outstanding 
seeing (~0.3 arcsec) in the night logs, and Mark said that this could be 
the result of new collimation procedures being used.


 	      3.5-m Telescope and Instruments Highlights
 	                09/25/14 through 10/22/14

0) Overview

Weather finally broke for the start of October and we resumed normal 
observing after being shut out for nearly all of September.

Gretchen Van Doren who has been with APO for 24 years is retiring at the 
end of the year.  A replacement for her has been hired and will start 
November 1.  Her name is Beth Mitchell.  Beth is new to APO and will take 
some time to settle in and even begin to fill the huge void that Gretchen 
will leave.

1) Telescope

The new TCC has been completed and transitioned to maintenance mode with a 
few minor upgrades still being pursued.  There have been a couple minor 
issues with tracking and learning the new features of the TCC.

2) Instruments

   o DIS has been operating normally;

   o Agile is operational;

   o TripleSpec had no operational problems this period;

   o SPIcam was operational during this period;

   o Echelle inter-order light is good after last months work with no 
additional issues;

   o NIC-FPS continues to operate in shared-risk observing mode due to 
occasional image corruption issue.  NFS less than 9 is operational;

   o GIFS is operational. GSFC personnel came to APO and after working on 
the dewar more than doubled the LN2 hold time.


The instrument development is proceeding on schedule.  The CCD we are 
purchasing did not meet specifications and so we will be testing and 
integrating using a loaner CCD until a new batch is made and we can get a 
good one.  The dewar integration with the engineering grade detector is 
nearly complete.  We will not be able to incorporate single quadrant 
readout as initially hoped due to the way the chip is wired.  We are 
finalizing the optics quote.


Georgia State joining ARC:

Suzanne reported that the ARC Board has approved the GSU induction into 
ARC.  Theirs will be a 1/16th share, beginning on 1 January.  Naturally 
this is cause for much celebration, and we welcome GSU as a new ARC 


ARCSAT update:

Mark mentioned that there has been recent engineering work done on ARCSAT, 
including the realuminization of the primary mirror.  Other improvements 
include upgrades to the telescope control system, thanks to help given by 
Jack Dembicky.  We are working on an option for manual focussing. The 
present system defaults to autofocus, which is not always needed.  The 
Quarter 4 schedule, i.e., mid-October through early February, has been 
issued and the observing has begun.  We are still operating in shared-risk 
mode.  In Quarter 1, we may start implementation of a system to charge 
users for telescope time.  There is also a tentative plan this quarter to 
trade ARCSAT time with UWisconsin for time on the WIYN 0.9-meter at Kitt 
Peak.  Whether this continues into the future is undetermined at present.
   Suzanne stressed that it is important that ARCSAT users always fill out
   the night log, which is the means for problem reporting, general user
   feedback, and how we derive user statistics.  Furthermore, ARCSAT users
   should log in to TUI at the beginning of each night and notify the 3.5-m
   Observing Specialist that they are going to be observing.


Old versions of TUI:

Although it is not a serious problem, Mark noted that some users are using 
2012 versions of TUI.  We encourage users to always use the current 
version of TUI, especially now that we are running under the new TCC.  An 
updated Windows version of TUI is pending.  The current released version 
of TUI (for Macs, linux) is 2.3.1.  The beta version, 2.3.2, is OK to use, 
especially if you need any of its new features that are self-documented in 
the version history.


NRC O/IR committee:

Rene reported on a recent meeting of the NSF/NRC Optical-Infrared 
committee that he participated in a couple weeks ago. He noted that LSST 
will not be doing their own follow-up observations so there is a lot of 
discussion and planning about what telescopes will be available.  There is 
renewed interest in resurrecting the AURA Coordinating Council of 
Observatory Research Directors (aka ACCORD), as well as exploring 
inter-observatory time trades, and robotic telescopes.  The future of WIYN 
has improved some with a new partner and possible NASA support.  The NSF 
support levels remain constrained.  The OIR study report (expected in a 
few months) will have specific recommendations that may impact APO.


APOGEE fiber link to 3.5m:

John Bally said that Colorado astronomers are enthusiastic about getting a 
fiber feed (single, or IFU) from the APOGEE NIR spectrograph at the Sloan 
Foundation Telescope run to the 3.5-m telescope.  There are exciting 
scientific possibilities for doing NIR multi-object spectroscopy of young 
stars.  There followed an extended discussion about establishing the fiber 
link that ranged from a simple single-fiber feed to a more complicated IFU 
system.  Suzanne said that we need a larger body of compelling scientific 
motivation for this project if it is to be implemented soon, especially 
since there will likely be significant scheduling issues while SDSS-IV is 
operating (until 2019-20).


Other business

Michael Strauss asked about the migration of the apo35-general mailing 
list from Princeton to APO.  Bill Ketzeback and Jon Brinkmann have the 
list set up, and we will implement it in January when Princeton and 
Chicago are no longer ARC members.  This will allow us to remove old users 
from the list during the switchover.  The 35-general email archive also 
needs to be migrated.



[open from previous meeting]:

=====> Action: Users committee representatives to gather science slides 
for Board report and send to Suzanne and Mark by October 15.

 	Status: A few institutions were reminded and promised to respond 
ASAP.  Action closed.

=====> Action: users committee to gather information from users about 
scientific interest in a fiber feed to APOGEE (2.5m) and smaller 
telescopes on site.

 	Status: Remains open for another month to gather additional 
scientific input.

[new from this meeting]:



Next meeting:  The next Users Committee phonecon will be on Monday 17 
November.  The agenda and other materials will be sent to the committee 
members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1346 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to